# CGCOD: Class-Guided Camouflaged Object Detection

### v1

@@ 6 Conclusion
In this paper, we propose a novel task, termed Class-Guided Camouflaged Object Detection, which employs class-level textual information to guide COD. To facilitate this task, we have meticulously constructed a large-scale dataset, CamoClass, which not only comprises real-world images under diverse and complex conditions, but also includes corresponding class text of objects. Building upon this foundation, we introduce a multi-stage reasoning framework, which include the class prompt generator (CPG) and the class-guided detector (CGD). Extensive experiments demonstrate the effectiveness of the proposed method and its superior performance compared with the existing COD method. Extensive experiments prove that our method has strong scalability in COD and can successfully detect objects of the unseen classes. Moreover, as a plug-and-play component, the CPG significantly enhances the performance of existing COD detectors. We hope that our exploration will inspire future research, for instance, investigating multi-object camouflage segmentation and examining whether COD models truly capture the semantics of camouflage.

### v2

@@ 6 Conclusion
To overcome the limitations of the existing COD paradigm that rely solely on visual features, in this paper, we propose a novel task, termed Class-Guided Camouflaged Object Detection (CGCOD), which employs class-level textual information to guide COD. To facilitate this task, we meticulously construct a large-scale dataset, CamoClass and propose a novel framework CGNet for CGCOD, which include the class prompt generator (CPG) and the class-guided detector . Comprehensive experiments validate the effectiveness of the proposed CGNet, demonstrating superior performance compared to existing COD methods. Additionally, the CPG acts as a plug-and-play component, significantly enhancing the performance of current COD detectors. We aim for our CGCOD research to provide valuable insights into camouflaged object perception and inspire more deep COD studies incorporating multi-modal auxiliary information.